### 19 Dec 2023 v1

# An Adaptive Placement and Parallelism Framework for Accelerating RLHF Training

Abstract
Recently, ChatGPT or InstructGPT like large language models (LLM) has made a significant impact in the AI world. These models are incredibly versatile, capable of performing language tasks on-par or even exceeding the capabilities of human experts. Many works have attempted to reproduce the complex InstructGPT’s RLHF (Reinforcement Learning with Human Feedback) training pipeline. However, the mainstream distributed RLHF training methods typically adopt a fixed model placement strategy, referred to as the Flattening strategy. This strategy treats all four models involved in RLHF as a single entity and places them on all devices, regardless of their differences. Unfortunately, this strategy exacerbates the generation bottlenecks in the RLHF training and degrades the overall training efficiency. To address these issues, we propose an adaptive model placement framework that offers two flexible model placement strategies. These strategies allow for the agile allocation of models across devices in a fine-grained manner. The Interleaving strategy helps reduce memory redundancy and communication costs during RLHF training. On the other hand, the Separation strategy improves the throughput of model training by separating the training and generation stages of the RLHF pipeline. Notably, this framework seamlessly integrates with other mainstream techniques for acceleration and enables automatic hyperparameter search. Extensive experiments have demonstrated that our Interleaving and Separation strategies can achieve notable improvements up to 11$\times$, compared to the current state-of-the-art (SOTA) approaches. These experiments encompassed a wide range of training scenarios, involving models of varying sizes and devices of different scales. The results highlight the effectiveness and superiority of our approaches in accelerating the training of distributed RLHF.
Index Terms:
Distributed Training, Large Language Model, RLHF, Heterogeneous Hardware
I Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across various natural language processing tasks[2, 6, 36, 42]. However, it is crucial to fine-tune LLM models to align them with human preferences. In the absence of proper alignment with human feedback, LLMs may exhibit behaviors that deviate from expectations. These behaviors include generating fabricated information, pursuing inaccurate objectives, and producing harmful, misleading, and biased expressions[25, 17, 12]). To tackle this issue, researchers have proposed a series of approaches aimed at ensuring that LLMs adhere to ethical considerations and societal norms.
One of the most successful approaches in addressing the alignment issue is Reinforcement Learning from Human Feedback (RLHF)[32, 1, 25]. Particularly, the pipeline proposed in InstructGPT [25], a sibling model of ChatGPT, has gained significant attention. RLHF combines human alignment with Reinforcement Learning (RL), a decision-making approach that allows AI agents to learn from interactions with the environment, receiving feedback in the form of rewards. This integration of RL with Human Feedback (HF) enables researchers to enhance the capabilities of LLMs and improve their alignment with human values. RLHF represents a significant advancement in building more reliable and responsible LLMs.
The RLHF training pipeline[25] typically follows a systematic procedure involving three steps:
[…]
The mainstream RL algorithm to finetune LLM is Proximal Policy Optimization (PPO) [28], as applied in InstructGPT. It leverages an Actor-Critic architecture[22], comprising four interconnected models: a Reference(Ref) model, a Reward model, an Actor model, and a Critic model. The Reference and Actor models are typically initialized from the SFT model, while the Reward model and Critic model are initialized from the Reward Model in Step 2 of the process. In the generation stage, the Actor model generates the responses from the prompts. Subsequently, given the prompts and response, all four models generate or infer intermediate output results such as logits or values. These results are stored in the experience replay buffer similar to DQN[23]. During the training stage, the Actor and Critic models are trained using batch data sampled from the buffer. This whole procedure repeats over a number of steps until converges.
Across various training scenarios, the size of individual LM parameters can exhibit significant variations, e.g. spanning from 7 billion to 65 billion parameters in models like the Llama series [33]. Furthermore, the presence of four LMs in RLHF pipeline amplifies memory usage quadratically, which underscores the critical need for efficient distributed training methodologies. However, traditional distributed training techniques, such as data parallelism or model parallelism[14, 39, 27, 31], are primarily designed for training a single model, not for efficiently managing multiple interdependent models as seen in RLHF. This necessitates the development of innovative model placement strategies. Existing works[40, 3, 34] typically employ a straightforward model placement strategy, namely Flattening strategy. This strategy treats all involved models as a single entity and applies standard parallelism techniques designed for individual models. Also, these frameworks follow a serial execution as the standalone PPO algorithm. However, this approach leads to fixed parallelism across all models, resulting in reduced memory available for a single model on a single machine. Therefore, we have to increase the parallelism level to accommodate all four models, resulting in additional communication costs [27, 38]. Furthermore, different scales of model sizes or devices may require distinct model placement strategies. For instance, a single device could accommodate smaller models with 7 billion parameters without introducing additional parallelism, unlike larger models with 65 billion parameters. This discrepancy results in memory redundancy and extra communication costs under a fixed model placement strategy. Lastly, the training and inference stages of the models inherently differ, necessitating distinct distributed strategies. Specifically, in the training stage, it is advisable to shard the optimizer state, gradients, and parameters sequentially for better efficiency. On the other hand, during the generation stage, only the parameters need to be shared, as they are the only components utilized during inference.
Therefore, we propose the Adpative Placement and Parallelism (APP) framework, which could significantly enhance the training efficiency in distributed RLHF training while ensuring no compromise on model performance. Firstly, our framework introduces two innovative model placement strategies, including Interleaving, and Separation strategies. They are designed to improve training efficiency in various RLHF training scenarios. Secondly, we design an APP Execution Engine abstraction that provides a user-friendly interface to modeling experts, shielding them from the intricate details of model placement and parallelism strategies. Lastly, we propose an Adaptive Parallel Planner that automates the complex model placement configurations, selecting the most suitable distributed model placement strategies for optimal performance. Furthermore, we are actively working towards releasing the code for the APP framework as an open-source project, pending approval from the company.
Our work makes four main contributions:
1. Adaptive Model Placement Strategy: Our APP framework introduces two flexible model placement strategies. The Interleaving strategy efficiently reduces memory redundancy and minimizes the number of communication participants in small-scale training by decoupling independent models. The Separation strategy further rebalances the training and inference workloads by decoupling the training and generation stages in large-scale training scenarios. This separation allows for the utilization of various inference optimization techniques, further enhancing the training efficiency.
2. Compatibility and Usability: Our APP framework incorporates an APP Execution Engine, which provides support for various model placement strategies and popular acceleration techniques while requiring minimal code modifications or even no modifications at all. The Adaptive Parallel Planner in the APP framework further simplifies the distributed configuration by automating the complicated model placement. This empowers users without distributed training knowledge to easily leverage the benefits of distributed RLHF training.
3. Heterogeneous GPUs Support: The APP framework takes into account the heterogeneity of mixed GPU capabilities through our Separation strategy. This enables the training of RLHF models, traditionally performed on high-end GPU devices, to be conducted on a broader range of more affordable and heterogeneous GPU clusters.
4. Superior Performance: The extensive experiments confirm that our approach could outperform the state-of-the-art solution by up to $\times1$, showcasing superior performance in terms of throughput, memory utilization, and overall training efficiency. Importantly, these improvements are achieved without any compromise on the effectiveness of the trained models.
II Background
II-A Data and Model parallelism
According to different parallel objects, distributed training technology for parallelizing single language model can be divided into data parallelism and model parallelism.
Data Parallelism (DP) involves dividing the input data into multiple shards and assigning them to different devices for parallel processing. In traditional DP training methods like AllReduce[11, 29] or Distributed Data Parallel (DDP)[20, 26], each device holds a complete replica of the model parameters and uses the AllReduce primitive to synchronize the model parameters across devices. ZeRO (Zero Redundancy Optimizer)[27] is a technique that addresses memory redundancy in DP training, also known as ZeRO-0. It categorizes the model states into three parts: optimizer states, gradients, and parameters. Notably, in mainstream mixed precision training with the Adam optimizer[18, 27], the memory consumption of optimizer states is roughly six times larger than that of parameters or gradients. To optimize memory redundancy, ZeRO performs sequential optimization in three stages: ZeRO-1, ZeRO-2, and ZeRO-3. In ZeRO-1, only the optimizer states are sharded across devices. In ZeRO-2, the gradients are additionally sharded, introducing additional reduce-scatter communication after each backward stage. The increased level of ZeRO parallelism helps reduce memory redundancy but comes with increased communication costs[27, 38].
Model Parallelism, including Tensor and Pipeline Parallelism, involves distributing model parameters across multiple devices within or across instances[31]. Tensor Parallelism (TP)[30, 31] achieves this by vertically splitting the model, sharding tensors across multiple devices, and computing through distributed matrix multiplication. However, due to the frequent global communication, Tensor Parallelism is more efficient within a single node with a high GPU interconnect bandwidth. As the number of nodes increases, its scalability becomes limited. On the other hand, Pipeline Parallelism (PP)[14, 10, 13] adopts a horizontal model splitting approach, assigning different layers to different devices. It usually uses micro-batching to hide the bubbles which is the idle duration of the device in the pipeline.
II-B Heterogeneous Network
Clusters with heterogeneous networks are common in the majority of cloud computing environments[16]. In high-performance clusters like NVIDIA DGX-2 or DGX-A100, these devices are equipped with NVLink/NVSwitch as intra-node interconnects, providing up to 4.8TGbps bandwidth. However, the bandwidth between inter-nodes, whether InfiniBand or Ethernet, is only 200-800Gbps. This means that the intra-node bandwidth can be approximately 24 times faster than inter-node bandwidth[38, 43, 35]. This limitation becomes even more pronounced in heterogeneous GPU clusters, such as a mixture of A100 and V100, where different types of GPU devices can only be connected using Ethernet. The significant disparity in bandwidth within and between nodes greatly hampers the communication efficiency of Tensor Parallelism and ZeRO, especially at higher levels of ZeRO[38, 27]. ZeRO-3, for instance, requires frequent AllGather communication primitives across all nodes to collect model parameters which are shared across all devices, as compared to ZeRO-2. Meanwhile, tensor parallelism requires more global communications for each matrix computation during computation. Moreover, as the size of the cluster scales out, the communication overhead caused by heterogeneous networks grows larger.
III Related Works
Since the emergence of ChatGPT and InstructGPT, several distributed frameworks have been proposed to support the parallelization of complex InstructGPT-like RLHF training pipelines. These works include trlX [3], Colossalai [21], Deepspeed-Chat [40], and trl [34], all of which employ the Proximal Policy Optimization (PPO) algorithm for RLHF training. Unlike standard approaches for training a single model, distributed RLHF training involves managing the model placement and computation of multiple models that are involved in the training process.
These works primarily consider two RLHF model structures: AC-Share[3] and AC-NonShare[40], depending on whether the Actor and Critic model share the parameters or not. For example, in the AC-Share structure, the Actor and Critic models share parameters by adding an extra linear layer to differentiate between them. This reduces the memory and resource requirements for RLHF training, but it may potentially lead to a loss in model performance.
It is important to note that all of these works adopt a fixed model placement strategy or its variant, referred to as the Flattening strategy, to parallelize the computation procedure of multiple models in the RLHF pipeline[40, 34, 21, 3]. The Flattening strategy straightforwardly treats all four models (Reference, Reward, Actor, and Critic) as a whole and distributes them across multiple acceleration devices. For instance, DeepSpeed-Chat[40], one of the popular RLHF training frameworks using the AC-NonShare structure, places all four models on each device and applies data and model parallelism techniques like ZeRO to parallelize the training. TrlX[3], another well-known work utilizing the AC-Share structure, employs a variant of the Flattening placement strategy. It places the Reward model on a fully occupied device to leverage its computational capabilities exclusively, while deploying the Actor, Critic (shared with the Actor model), and Reference model on other devices using the Flattening strategy.
However, there is an apparent issue in the trlX framework. By placing the Reward model on a single GPU device, it underutilizes both memory and computational power. If the Reward model size is too small, GPU memory is wasted, while a Reward model that is too large may lead to out-of-memory issues. Additionally, the device exclusively allocated to the Reward model will experience significant idle time within a single step. Furthermore, extra communication is required to gather the generated results using AllGather from the Actor model, and Scatter the prediction outputs back to the Actor model placed on other devices. These problems significantly degrade the overall efficiency.
IV Limitations and Challenges
Figure 1: The Percentage of Each Stage Spent in the RLHF Training Using DeepSpeed-Chat.
IV-A Limitations in Existing Distributed RLHF Training
Currently, the generation stage significantly slows down the entire RLHF training. Without loss of generality, we take the AC-Share structure as an example. Figure 1 illustrates that the Actor Generation model accounts for more than 85% of the total duration in one step, while the Training stage only occupies 10% and then other forward stages using DeepSpeed-Chat. The bottlenecks come from two main reasons: Sequential Execution and the fixed model placement strategy, namely the Flattening strategy.
[…]
Firstly, the Flattening model placement strategy incurs memory redundancy and additional communication costs in the generation phase. It deploys all four prediction models on every device, significantly reducing the available memory on each device for a single model. As a result, parallelism techniques like ZeRO are employed to handle these large models, but they come with the drawback of memory redundancy or communication costs. For instance, in deepspeed-chat with ZeRO-1, redundant gradients, and model parameters are maintained for all models on every device. Also, the communication cost rockets up with the increased level of ZeRO parallelism as discussed in sectionII. This problem is significantly exacerbated in RLHF training scenarios where dozens or even hundreds of GPU devices are typically involved. In such cases, the memory redundancy and communication costs are directly proportional to the scale of the computing devices, exacerbating the problem.
Secondly, the training and inference stages naturally differ from each other, while the Flattening strategy fails to decouple them, especially for the Actor and Critic model, resulting in low efficiency. In terms of parallelism, the training phase requires taking all the model states including parameters, gradients, and optimizer states into consideration, while only parameters are used in the inference model, which requires different parallelism techniques. For example, Tensor Parallelism or PaRO[31, 38] could be used to parallelize the model inference rather than ZeRO parallelism which fits model training. These conflicts of parallelism between training and generation stages lead to efficiency degradation. Additionally, these two stages require different optimization techniques, such as PagedAttention in vLLM[19], quantization[7, 41] or speculative sampling[4] which are used to optimize generation or inference workloads.
Lastly, the usage of heterogeneous devices in a GPU cluster is hard to achieve under such a strategy. As mentioned earlier, RLHF training typically involves dozens, or even hundreds, of GPU devices. However, even in industry, scheduling hundreds of homogeneous high-end GPUs often results in long queuing duration. On the other hand, obtaining heterogeneous GPUs (e.g., a combination of A100 and V100) is comparatively easier [15, 37]. However, training tasks with such heterogeneous GPUs are significantly more difficult than inference tasks, mainly due to kernel compatibility and network heterogeneity[8]. The Flattening strategy treats both the generation and training models as a whole, making it difficult to leverage the heterogeneous devices for simpler inference workloads in RLHF training pipeline.
IV-B Key Challenges
To overcome these limitations, it is crucial to explore new distributed model placement strategies for efficient RLHF training, which also could save substantial costs for large-scale training prevalent in this scenario. However, to build an efficient and easy-to-use RLHF training framework, we encounter multiple challenges, including:
[…]
Figure 2: Our APP Framework
V Our Framework
Therefore, we introduce our Adaptive Placement and Parallelism (APP) Framework for accelerating the RLHF Training. The APP framework carefully orchestrates the RLHF pipeline while striking a balance between memory utilization and communication in distributed computing. Moreover, it caters to different training scenarios, encompassing varying model scales and heterogeneous devices. Our APP Framework consists of three components: Model Placement Strategies, APP Execution Engine and Adaptive Parallel Planner, as illustrated in Figure 2.
Firstly, besides the Flattening strategy, we introduce two model placement strategies aimed at enhancing the efficiency of RLHF training. These strategies are carefully designed to optimize communication efficiency and effectively utilize memory, catering to different training scenarios that involve different model scales and hardware heterogeneity. Secondly, we present the design of an APP Execution Engine abstraction, which provides modeling experts with a simplified programming API by decoupling RLHF modeling from distributed computing. Lastly, we extend our work by developing an Adaptive Parallel Planner, capable of automatically configuring the distributed related hyperparameters to achieve maximum throughput during RLHF training at best effort.
Figure 3: The architecture of Model Placement Strategies, where (a) represents the Flattening strategy, (b) represents the Interleaving strategy, (c) demonstrates the Separation strategy used for homogeneous devices, where generation and training models are assigned to exclusive groups of devices, and (d) showcases the Separation strategy used for heterogeneous devices, where inference models are allocated to dedicated groups of devices specialized for inference.
V-A Model Placement Strategies
Figure 4: The timeline of Flattening strategy vs. Interleaving strategy. Two homogeneous GPU devices are utilized for demonstration. Under the Flattening strategy, both the Reference model and Reward model are deployed on both devices, whereas under the Interleaving strategy, the Reference model and Reward model are allocated to separate devices.
V-A1 Interleaving Strategy
The main idea of the Interleaving strategy is to partially parallelize the pipeline by allocating the independent models on exclusive group of GPU devices, as well as shrinking the number of computing devices required for each model. We observe that either Reward&Reference Inference (and Actor&Critic training models) could operate independently, thereby it is possible to eliminate part of the serial execution. For example, the Reference model and Reward model both take the generated responses from the Actor model as input and output the intermediate results into the experience buffer without synchronization. Also, the Actor and Critic do not need synchronization duration of the training phase. This design helps the parallelization of independent models in time-cost RLHF training. This lowers the memory redundancy and shrinks the scale of participating devices used by either the Reference or Reward model rather than using all devices in the Flattening strategy.
We compare the distributed architecture and computation procedure between the Flattening strategy and our Interleaving strategy, using a small number of devices as illustrative examples. Our strategy could be easily extended to large-scale computing devices. In terms of Model Placement, the Flattening strategy treats the four models as a unit and deploys them on all devices, as depicted in Figure  3(a). By contrast, our Interleaving strategy distinctly assigns the Reference model and Reward model to two distinct device groups (GPU 0-3 and 4-7) [^1] Our interleaving strategy could further place the Actor and Critic model on exclusive devices. We only illustrate the interleaving between Reward and Reference for simplicity., as illustrated in Figure  3(b). Our Interleaving placement strategy reduces the number of devices used by either the Reference model or the Reward model. It’s important to note that responses generated from the Actor are located on GPUs where the Reference model is not placed since the Actor model is deployed on all GPU devices, as shown in the timeline in Figure  4(b). Therefore, it is necessary to perform two extra communication stages: an AllGather primitive to gather the generated results from all devices before forwarding, and then perform a Scatter primitive to distribute the results back to all devices. However, these two additional communication stages can be considered negligible due to overlapping computation and communication. We conclude that the Interleaving strategy shortens the time-cost generation duration from two aspects.
Firstly, in terms of memory consumption, our Interleaving strategy reduces the number of participating devices for either the Reference or Reward model by half, effectively minimizing memory redundancy, especially in ZeRO-0 parallelism. This allows our method to leverage the saved memory to increase the total batch sizes, resulting in higher overall throughput. For instance, when employing the Flattening strategy under ZeRO-0, all devices hold replicas of the model states, while only half of devices need to hold the replicas under our Interleaving strategy. As depicted in Figure  4(b), our method can accommodate a larger batch size compared to the Flattening strategy, as shown in Figure  4(a).
Figure 5: The timeline of Separation strategy on homogeneous (a) and heterogeneous GPU devices (b). In the homogeneous Separation strategy, the training and inference models are placed on GPU 1&2 and GPU 3&4 respectively. In the heterogeneous Separation strategy, these two stages are allocated to dedicated devices exclusively for either training or inference purposes.
Secondly, from a communication efficiency perspective, having few devices leads to reduced communication overhead in both the generation and training stages. The communication cost increases directly with the number of participating devices, particularly for inter-node communication. It’s important to note that while increased levels of ZeRO parallelism reduce memory redundancy in model states, it comes at the cost of increased communication[27, 38]. For instance, ZeRO-3 Parallelism allows for accommodating larger models in less memory. However, during both the generation and training stages, an AllGather operation is required to obtain all parameters of the current layer before usage. As the scale of computing nodes increases, the problem is exacerbated, especially in GPU clusters with low-bandwidth inter-node communication, which is the bottleneck in current distributed LLM training. However, the Interleaving strategy can further leverage high-speed intra-node bandwidth and reduce inter-node communication by colocating the Reward model or Reference model on a single machine with usually 8 GPU devices, particularly in small to medium training scenarios. This approach significantly decreases communication overhead, especially inter-node communication, and improves overall efficiency.
Furthermore, we eliminate the Forward step of the Actor model by aggregating it into the Actor generation step, as indicated by the dashed box in Figure  4(a). This aggregation in computation steps also brings certain efficiency gains.
V-A2 Separation Strategy
To shorten the generation duration, our Separation strategy involves separating the training stage and the generation stage of the RLHF pipeline. We recognize that the training and inference stages are inherently different. This division does not only allow us to run the training and generation models simultaneously but also leverage various powerful inference optimization techniques and idle heterogeneous devices in clusters. It could largely accelerate the generation stage and re-balance the workloads between these two stages in RLHF pipeline.
This strategy significantly speeds up the entire RLHF training, although it introduces certain memory redundancy and communication overhead. Specifically, in our strategy, the Actor and Critic models are classified into Prediction(Pred) Actor/Critic models and Training Actor/Critic models by duplicating, which are exclusively used for training or inference. Figure  4(c) illustrates the placement of the training Actor and Critic models and other prediction models, including the Pred Actor and Critic models, on separate devices. Optionally, we could further place the training Actor and Critic models on exclusive devices in an interleaving manner to further decrease the memory redundancy and communication costs. However, there is a need for communication for intermediate results and parameter synchronization between the training Actor/Critic models and the prediction models. Firstly, as shown in Figure  5(a), extra point-to-point communications are required before or after the Forward stage of each prediction model. Additionally, the training Actor and Critic models need to pull the intermediate results from the inference devices. These extra communication stages introduce some idle time per device, known as bubble overhead. However, these bubbles can be minimized by overlapping computation and communication stages through the use of pipeline parallelism. As illustrated in Figure  5(a), the bubbles are nearly hidden through micro-batching in a pipeline parallelism manner, especially with a higher number of generation times. Secondly, it is necessary to synchronize the trained model parameters between the training Actor/Critic models and the inference models after each round of PPO training. The synchronization overhead is inevitable but acceptable.
Although this approach introduces some complexity and overhead, it provides significant benefits from three different perspectives:
Firstly, from the perspective of parallelism, our Separation strategy decouples the training and generation stages for the Actor/Critic models, as discussed in Section IV. This allows us to apply suitable parallelism strategies to either training or the Pred Actor/Critic model. For example, the time-consuming Pred Actor model can be accelerated by utilizing specialized parallelism strategies that are well-suited for inference workloads, such as intra-node tensor parallelism. To be more specific, the Pred Actor model can be replicated on different machines, and the parameters can be sharded horizontally or vertically across the devices within the same machine. This optimization fully takes advantage of the high-bandwidth intra-node communication and minimizes inter-node communication.
Secondly, the separation between the training and generation stages allows the APP framework to easily leverage the rapidly evolving inference optimization techniques by integrating different open-source libraries, such as quantization[7, 41] or speculative sampling[4]. In contrast, under the Flattening strategy, it is challenging to fully take advantage of these optimization techniques due to the usage of the same model for both training and prediction stages. Also, the frequent swapping between training and prediction modes of the same model incurs extra costs.
Lastly, this strategy allows the APP framework to scale out additional computing resources, including heterogeneous devices, for the time-consuming generation stage. This scalability helps rebalance the throughput between the training and generation stages, resulting in accelerated overall training efficiency. With this approach, high-end GPU devices like A100 can be used for training models, while more cost-effective GPU devices such as NVIDIA A10 or even 3090 can be utilized for inference models. This flexibility in either quantity or type of devices offers significant adaptability to diverse training scenarios.
V-B APP Execution Engine
The APP Execution Engine is designed to provide a user-friendly interface for modeling experts to easily build their own RLHF training pipeline. It is compatible with various optimization techniques and supports different model placement strategies to meet the requirements of various scale training scenarios. Our placement strategies may involve 3 to 6 models, including Pred Actor, Pred Critic, Reference, Reward, Train-Actor, and Train-Critic, along with heterogeneous devices based on the chosen placement strategy as well as AC share or AC Non-Share structure.
Model Placement Abstraction. The APP Execution Engine encapsulates each model involved in RLHF pipeline considering both AC share and AC Non-Share structures. Our modular design allows for the encapsulation of each model and corresponding distributed techniques. These techniques include the model placement strategy, singular model parallelism (such as ZeRO in DeepSpeed and Tensor Parallelism in Megatron), as well as various optimization techniques in other libraries. While our placement strategies offer general guidelines on how to distribute these models, there remains a gap in determining the optimal number of devices for deploying a target model, as this relies on factors like model sizes and available hardware in different training scenarios. To address this, we introduce the concept of Model Placement Ratio, which is a value ranging from 0 to 1 that represents the proportion of devices allocated for a specific model. This provides enhanced adaptability. For instance, the Flattening strategy for a Reward model has a ratio of 1, indicating that the model is deployed on all devices. On the other hand, the Interleaving strategy for the Reward model, as depicted in Figure 4, has a ratio of 0.5, signifying that the model is deployed on half of the devices. Then the engine constructs a communication subgroup on the deployed devices of the Reward model for synchronization. It also belongs to a parent communication group that is connected with all models’ subgroups for data interaction. Then our Execution Engine generates the model placement mapping, which places the model to physical devices properly according to the model placement ratio, considering the device capacity and topology, such as intra- and inter-node. This mapping involves detailed ranking assignments for each model. Overall, this abstraction effectively shields users from the complexities associated with model placement, parallelism techniques, device computing capacity, and topology.
Data Interaction. Data interaction caused by the placement strategy is also encapsulated within the Execution Engine. The Interleaving and Separation strategies introduce communication among different communication subgroups during the different training stages, and the APP handles these data interaction logics. Each training or prediction operation of the models is treated as an individual Operator. The rank assigned to each model performs an AllGather operation to collect the upstream data before executing the Operator. Once the Operator is executed, the resulting data is scattered back downstream.
This abstraction enables a general interface and backend for creating a wide range of RLHF algorithms for research or industrial applications. Users can focus on the essence of RLHF modeling without being burdened by the intricacies of the underlying implementations.
V-C Adaptive Parallel Planner
Figure 6: The Flowchart of Adaptive Parallel Planner
The objective of the Adaptive Parallel Planner is to search for hyperparameters, including optimal model placement mapping, to maximize the training throughput of RLHF in various training scenarios while considering computing resource and model size constraints. Once the hyperparameters are determined according to the maximum overall floating-point operations per second (TFLOPS), the APP Execution Engines utilize it to execute the RLHF pipeline finally.
The workflow of the Adaptive Parallel Planner is illustrated in Figure 6. The input of the planner consists of NodeDesc (specifying the quantity, type, and memory of GPU devices), RLHFModelDesc (indicating the size of each model), and an optional Model Placement strategy which could be represented by Model Placement Ratio. The output includes hyperparameters, with the key model placement ratio being a crucial component. The hyperparameters also include batch size during the generation and training stages, as well as the Parallelism method (like ZeRO 1 or 2).
The workflow can be summarized as follows:
1. The MinResourceChecker initially calculates the minimum GPU resources required based on the NodeDesc, RLHFModelDesc, and optional Model Placement Ratio. If the available computing resources do not meet the minimum requirement, the process is halted.
2. Subsequently, the Planner explores the best configurations by using two search strategies. The first is a prior-based strategy that searches for the optimal configuration according to predefined rules. This strategy considers our prior knowledge, such as allocating resources ranging from 30% to 50% for Pred Actor models and training models respectively, while reserving the remaining resources for other prediction models. The second strategy is a grid search-based approach that exhaustively explores different combinations.
3. The Planner begins by iterating through a set of candidate model placement ratios for all four models. Each candidate placement ratio is represented as a vector, with each element corresponding to the model placement ratio. Next, the Planner proceeds to iterate through candidate hyperparameters for each model involved in the RLHF pipeline, including both the training and inference stages. In this step, NVIDIA NeMo[24] can be utilized to search for hyperparameters specific to each model. Once the hyperparameters for the RLHF pipeline are determined, they are passed on to the APP Execution Engine. This engine runs the training for a specific duration and calculates the TFLOPS.
4. Lastly, the APP engine outputs the resulting memory utilization and TFLOPS measurements to the Judger. It determines whether to continue the hyperparameter search based on predefined conditions, such as the specified iteration or duration of searching. The hyperparameters that yield the highest TFLOPS are then utilized in the production RLHF training process.
VI Experiments
We conduct comparative experiments on both AC-NonShare and AC-Share model structures against the Flattening strategy used by DeepSpeed-Chat and trlX respectively. We evaluate the model throughput while using the maximum batch size available for each framework before Out-of-Memory (OOM) for fairness. These experiments encompass variations in model sizes, device scales, and device types, allowing us to assess the effectiveness of our APP framework across different training scenarios.
VI-A Experiment Settings
LLM Backbones. We employ Llama[33] of different parameter sizes as our LLM backbone, including 7B, 13B, 33B, and 65B. For demonstration purposes, we keep consistent model size for each model involved in RLHF pipeline, but our APP framework could adapt to training scenarios where each RLHF model may have varying model sizes and other LLM models, such as Opt[42] and GLM[9].
Hyperparameters. We have established the following parameters for our research. Both the number of PPO training epochs and batches per rollout are set to 1. The input sequence length is limited to 256. Both minimum and maximum generation lengths are 256. The batch size for each experiment is set to the maximum value before Out of Memory. For other model convergence-related parameters, we keep consistent for fairness.
Dataset. We perform RLHF on the default dataset for DeepSpeed-Chat. At the time of writing, “Dahoas/rm-static” hosted on HuggingFace is employed for tuning LLM via RLHF. This is an open-source ChatBot or Assistant style dataset, specifically designed to create a Helpful & Harmless conversational system[1].
Baseline. We conduct a comparison between the APP framework and the current state-of-the-art methods in RLHF training: DeepSpeed-Chat with Hybrid Engine for the AC-NonShare case and trlX for the AC-Share case.
Evaluations. We evaluate the effectiveness of different training frameworks by comparing their sample throughput during training. Sample throughput refers to the rate at which samples are processed end-to-end, typically measured in samples per second(# samples/sec). Furthermore, we verify that the model convergence, measured in returns, remains unaffected by our placement strategies, as evidenced in Figure 9, when compared to the DeepSpeed-Chat implementation.
Environments. Our experiments use the following software versions: CUDA 11.7, DeepSpeed 0.11.2, trlX 0.6.0, PyTorch 1.9.2, Megatron 3.0.2 and NCCL 2.14.3. The experimental cluster utilized in our study comprises up to 16 DGX nodes, with each node equipped with 8 Ampere A100 SXM3 80GB GPUs. The GPUs within each node are interconnected using NVLink, providing a high-speed bidirectional bandwidth of up to 600GB/s. In addition, the nodes are connected via 8 InfiniBand (IB) adapters, which support NVIDIA SHARP and provide around 100GB/s inter-node bandwidth.
VI-B Experiments and Analysis
We compare the throughput of our APP framework against DeepSpeed-Chat and trlX respectively. Specifically, in the training of the 7B model, we utilize ZeRO-2 for training models and ZeRO-0 for inference models. For the training of 13B, 33B, and 65B models, we employ ZeRO-3 for both stages. Furthermore, our Separation strategy incorporates tensor parallelism.
VI-B1 Comparing with DeepSpeed-Chat
Interleaving vs Flattening Strategy. We implement two Interleaving strategies to compare them with the Flattening strategy used in DeepSpeed-Chat. In $\text{Interleaving}_{1}$, the Reference and Reward models are placed on half of the GPU devices. In $\text{Interleaving}_{2}$, we interleave not only the Ref and Reward models but also the Actor and Critic models. Based on our experiments as illustrated in Table I, we conclude that the $\text{Interleaving}_{1}$ strategy is 12% faster than the Flattening strategy in DeepSpeed-Chat in small training scenarios, such as with 7B and 13B models. Additionally, the $\text{Interleaving}_{2}$ strategy is 71% faster than DeepSpeed-Chat in larger training scenarios, specifically with 65B models when employing ZeRO-3. The speedup ratio further increases when scaling out to more devices.
i) $\text{Interleaving}_{1}$ vs. Flattening Strategy. During the training of 7B size models using 1$\times$8 devices, we observe that $\text{Interleaving}_{1}$ strategy in our APP framework leads to a throughput increase from 16.12 to 17.35 samples per second, representing a speedup of 7.63% compared with Flattening strategy in DeepSpeed-Chat. This speedup ratio further improves to 12% when scaling out from 1$\times$8 to 4$\times$8 devices. The reason behind this improvement is that the adoption of the interleaving strategy effectively halved the memory redundancy of the Ref and Reward models. Consequently, the freed-up GPU memory can be utilized to expand the overall batch size, as explained in Section V-A. In this case, the Interleaving strategy achieves a maximum batch size of 36, whereas DeepSpeed-Chat’s maximum batch size is 24. However, it is important to note that $\text{Interleaving}_{1}$ does not perform well with larger model scales, such as in the 33B and 65B training scenarios. This is because these models require ZeRO-3 to reduce the memory redundancy to enable training large language models, while there is no memory redundancy of model states in this case.
ii) $\text{Interleaving}_{2}$ vs. Flattening Strategy. In contrast, the $\text{Interleaving}_{2}$ strategy compensates for this drawback in medium to large training scenarios. Although the $\text{Interleaving}_{2}$ strategy only allocates half of the original GPU count for the Actor model, it exhibits significant improvements when combined with the ZeRO-3 mode. As shown in Table I, the $\text{Interleaving}_{2}$ strategy is respectively 52% and 71% faster than DeepSpeed-Chat in the 33B and 65B model training scenarios when using 2$\times$8 devices. The reason behind this improvement is that, under $\text{Interleaving}_{2}$, the Actor and Reward models are placed exclusively on a single node, thus avoiding frequent inter-node AllGather communication primitives in both the prediction and training stages under ZeRO-3. Additionally, the number of communication participants for the Actor/Critic models is reduced by half.
TABLE I: Comparing Interleaving Strategy with Flattening Strategy in DeepSpeed-Chat. The results in the table indicate the throughput of the model, measured in samples per second (# samples/sec). The best results are highlighted in bold.
# of Strategy Model Size
GPUs 7B 13B 33B 65B
8 DeepSpeed-Chat 16.12 7.32 0.69 OOM
$\text{Interleave}_{1}$ 17.35 7.63 0.69 OOM
$2\times 8$ DeepSpeed-Chat 32.15 15.35 0.64 0.14
$\text{Interleave}_{1}$ 34.70 15.66 0.64 0.14
$\text{Interleave}_{2}$ 18.27 8.27 0.97 0.24
$4\times 8$ DeepSpeed-Chat 61.95 28.60 0.93 0.20
$\text{Interleave}_{1}$ 69.44 28.54 0.95 0.20
Separation vs Flattening Strategy. The experiments for the Separation strategy are specifically conducted in large-scale training scenarios, such as training 33B and 65B models on $4\times 8$ and $16\times 8$ GPU devices. However, this strategy is not well-suited for small training scenarios with a limited number of GPUs and smaller models. The reason is that the Separation strategy involves adding a replication of the Actor and Critic models exclusively for the prediction stage, which results in increased memory redundancy.
However, the benefits of the Separation strategy are significant, as demonstrated in Table II. Our Separation strategy outperforms the Flattening strategy used in DeepSpeed-Chat by factors ranging from 4$\times$ to 11$\times$. For instance, in 33B model training using 4$\times$8 GPU devices, the throughput increases from 0.93 to 4.76 samples per second, resulting in a speedup ratio of 4$\times$. The speedup ratio further increases to 7$\times$ when the model size is increased to 65B. Additionally, the Separation strategy achieves even higher speedup when scaling out to $16\times 8$ GPU devices, reaching a speedup ratio of 11$\times$ compared to the baseline using DeepSpeed-Chat. The contribution comes from that the Separation strategy decouples the training and generation stages of Actor model, which is not possible for the Flattening strategy. This decoupling allows the usage of intra-node tensor parallelism for the generation stages as discussed in section V-A. For example, in our training of 65B model using $16\times 8$ GPUs, $3\times 8$ devices are exclusively allocated for the Pred Actor model. In this case, we apply tensor parallelism within the same node while replicating it across three machines. This approach significantly reduces inter-node communications compared to frequent inter-node AllGather and Scatter communications in ZeRO-3, which are required in the 33B and 65B model training for the Flattening strategy.
TABLE II: Homogeneous Separation strategy vs. DeepSpeed-Chat.
# of Strategy Model Size
GPUs 33B 65B
$4\times 8$ DeepSpeed-Chat 0.93 0.20
Separation 4.76 1.75
$16\times 8$ DeepSpeed-Chat 2.11 0.55
Separation 10.07 6.80
TABLE III: Heterogeneous Separation strategy vs. DeepSpeed-Chat. We incorporate an additional $2\times 4$ V100 GPUs for inference models to speed up the generation stage.
# of Strategy Model Size
GPUs 33B
[…]
Heterogeneous Separation vs DeepSpeed-Chat. We also evaluate the Heterogeneous Separation strategy in a heterogeneous GPU cluster. The cluster consists of 8 A100 GPUs and an additional 8 spare V100 GPUs with 32GB memory. These heterogeneous V100 GPUs can not be utilized under the Flattening strategy. We specifically conduct experiments using large models, such as the 33B model. Under our heterogeneous Separation strategy, the Reference and Reward models are placed on a separate node with 4 V100 GPUs respectively. We enable the intra-node tensor parallelism for both models and allocate two A100 GPUs to two Pred Actor replicas exclusively. This configuration aims to accelerate the generation speed and improve the overall training speed.
As shown in Table III, introducing the idle V100 devices under the heterogeneous Separation strategy results in a speedup of 14.49%. This acceleration increases the throughput from 0.69 to 0.79 samples per second. Although the acceleration may not be considerable, this strategy explores a novel solution that allows RLHF training to utilize idle heterogeneous resources to enhance the overall throughput. It re-balances the inference workloads among different GPU devices by simply adjusting the batch size, without requiring extensive consideration of software and hardware heterogeneity among different types of GPUs.
VI-B2 Comparing with trlX
In this subsection, we compare our APP framework against trlX, which employs a variant of the Flattening strategy, in the AC-Share scenario. We evaluate our implementation of Flattening and Interleaving strategies in the APP framework in terms of throughput under various computing resources. However, before presenting the experimental results, there are three important notes to mention:
i) To align the number of models with trlX for a fair comparison, our Flattening and Interleave strategies only consider the Actor, Reference, and Reward models, excluding the Critic model in terms of placement. The reason is that the Actor and Critic models share most of the parameters in AC-share structure. The modifications required are minimal, thanks to the modular design of our model placement abstraction.
ii) We do not include comparative experiments involving the Separation or Heterogeneous Separation strategies because they are primarily effective for models with large parameter sizes, such as the 33B and 65B models. However, trlX can only place the Reward model on a single GPU, requiring the use of offloading techniques[26] to run the 33B and 65B models. However, offloading will significantly reduce the model’s prediction speed. Hence, it can be inferred that our Separation strategy offers even greater advantages in the trlX scenario compared to experiment results in DeepSpeed-Chat.
iii) In the trlX approach, there are two ways to allocate the Reward Model: standalone and coexisted. In the standalone mode, the Reward Model exclusively occupies a single GPU, while in the coexisted mode, the Actor model is evenly distributed across all GPUs, and the Reward Model is placed on one of them. We conduct a comprehensive comparison of these two modes to evaluate their performance.
TABLE IV: Comparing our Flattening strategy with trlX on $1$ machine with $8$ GPU devices. Two kinds of trlX implementation are employed.
Gradient Strategy Model Size
Checkpointing 7B 13B
OFF trlX (stand-alone RM) 6.04 2.84
trlX (coexisted RM) 6.20 2.20
Flattening 10.48 2.9
ON trlX (stand-alone RM) 5.92 2.87
trlX (coexisted RM) 6.21 2.22
Flattening 25.58 11.53
TABLE V: Comparing our Flattening and Interleaving strategy with trlX on varying numbers of machines, each with $8$ GPU devices.
# of Strategy Model Size
GPUs 7B 13B
$1\times 8$ trlX (coexisted RM) 6.21 2.22
Flattening 25.58 11.53
Interleaving 27.27 11.76
$2\times 8$ trlX (coexisted RM) 7.76 3.23
Flattening 50.67 22.26
Interleaving 54.71 24.63
$4\times 8$ trlX (coexisted RM) 8.87 4.24
Flattening 102.13 46.81
Interleaving 109.18 46.16
Specifically, in our experiments comparing with trlX, we have the following observations:
Figure 7: Training convergence between our proposed strategies and DeepSpeed-Chat.
Figure 8: Time cost between Separation strategy and DeepSpeed-Chat in generation and training.
Figure 7: Training convergence between our proposed strategies and DeepSpeed-Chat.
Figure 8: Time cost between Separation strategy and DeepSpeed-Chat in generation and training.
Figure 9: Time cost between Separation strategy and trlX in generation and training.
Flattening vs trlX Our Flattening strategy can achieve a speedup ratio up to $\times3$ compared with the trlX (with or without gradient checkpointing enabled). Gradient checkpointing[5] refers to whether activation value recalculations are enabled during training. We conduct the experiments for the 7B and 13B models on a single machine with 8 GPUs.
As shown in Table IV, our Flattening strategy without gradient checkpointing accelerated the throughput from 6.04 to 10.48 samples per second, resulting in a speedup of 73.51% for training the 6B models. This improvement arises due to the single-device bottleneck in trlX, where one device is responsible for aggregating data from other devices, incurring significant communication costs and leaving other devices idle, as discussed in Section IV. Moreover, the remaining memory that is not utilized by the Reward model is wasted, limiting the overall batch size. Also, this design leads to a longer generation duration, leaving other devices idle and resulting in underutilized resources, regardless of whether trlX is in standalone or coexisted mode. It is noteworthy that trlX narrows the performance gap when the reward model can fully utilize the single device in 13B model training. Additionally, when gradient checkpointing is enabled, the performance gap increases significantly, resulting in a speedup ratio of 332%. This is because gradient checkpointing can reduce memory consumption and in turn, increase the overall training batch size. However, the single-device bottleneck in trlX disables this benefit, hindering the potential performance improvement from gradient checkpointing.
Interleaving vs trlX. Our Interleaving strategy demonstrates significant improvements in training speed compared to trlX, which achieves acceleration ranging from 3.4×3.4$\times$3.4 × to $\times1$. As shown in Table V, our Interleaving strategy increases the throughput from 6.21 to 27.27 samples per second, resulting in a speedup ratio of 3.4$\times$ compared to trlX for the 7B model training on 1$\times$8 GPUs. This speedup ratio further increases to 6.1$\times$ and 11$\times$ when scaling out to 2$\times$8 and 4$\times$8 GPU devices respectively. Additionally, our Interleaving strategy improves the throughput by 6.6%, 8.3%, and 7% respectively when compared to our Flattening strategy, as we scale out to more GPUs. The training with the 13B models exhibits a similar pattern. It is analyzed that the single-GPU bottleneck in the trlX approach significantly limited the throughput, even using the maximum batch size before running out of memory under different resource configurations for the 7B and 13B models. Furthermore, the interleaving of the Reward model and Reference model in our Interleaving strategy reduces memory redundancy and decreases communication costs, resulting in considerable acceleration.
VI-B3 Time-cost Analysis
We conduct a further analysis of the time duration for the training and generation stages. This analysis confirms that our placement strategy significantly improves training efficiency by substantially reducing the time-consuming generation stage.
i) In the AC-NonShare scenario, we experiment with the Llama 33B model on 4$\times$8 GPU devices. Under DeepSpeed-chat’s Flattening strategy, the generation stage occupies the majority of the total duration, accounting for 93.31% in one step, while the training stage only takes up 6.69% as illustrated in Figure 9. However, our Separation strategy reduces the overall duration in one step from around 400 seconds to 30 seconds by significantly reducing the time-consuming generation stage. The percentage of the generation stage is reduced from 93.31% to 56%, almost half of the total duration. This reduction is particularly significant. It is attributed to our Separation strategy, which allows for intra-node tensor parallelism and other inference-optimized techniques for the inference models.
ii) In the AC-Share scenario, we examine the time expenses for the Llama 7B model on 2$\times$8 GPU devices. As depicted in Figure.9, our separation approach can reduce the overall duration from around 72 seconds to 10 seconds by largely decreasing the generation stage. The percentage of generation duration drops from 97.86% to 34.83%. In the generation phase, the trlX approach suffers from the aforementioned single-device bottleneck. In contrast, our Separation strategy can speed up the generation stage using various optimization techniques, resulting in significant reductions in overall training time.
VII Conclusion
In this paper, we introduce a novel RLHF training framework that adaptively places multiple large language models in the RLHF pipeline according to different training scenarios in terms of model and device scales. Our two placement strategies, Interleaving and Separation, not only significantly enhance training efficiency but also offer adaptable solutions for different training scenarios. Additionally, our framework is easy to use and provides automatic hyperparameter tuning for better efficiency. In the future, we plan to further improve our framework by enhancing the Planner component, such as exploring advanced techniques beyond prior or grid-based methods.